OpenStudy (anonymous):

solve the equation 3x^2+x=-7

OpenStudy (anonymous):

what does ^ mean

OpenStudy (anonymous):

dpalnc I WANT TO ANSWER!!

OpenStudy (anonymous):

DONT!!!

OpenStudy (anonymous):

Just need to know what ^ means

OpenStudy (anonymous):

get it in standard form: 3x^2 + x + 7 = 0 all yours.,..:)

OpenStudy (anonymous):

idk how to do it at all

OpenStudy (anonymous):

omg what does ^ mean

OpenStudy (anonymous):

to the second power

OpenStudy (anonymous):

\[3x ^{2}+x+7\]

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

ok so 3 times 3 equals 9... 9x+x+7= 0

OpenStudy (anonymous):

you can't do it like that because it's over the x

OpenStudy (anonymous):

awe ok one sec I got this lol

OpenStudy (anonymous):

You getting it?

OpenStudy (anonymous):

Omfg no im going crazy lol its impossible!

OpenStudy (anonymous):

@dpaInc help!

OpenStudy (anonymous):

what math is this?

OpenStudy (anonymous):

Use the quadratic formula because it is prime and can't be factorized so a=3, b=1 and c=7

OpenStudy (anonymous):

Yeah I was going to do that then i got lost

OpenStudy (anonymous):

You should get x = -1/6 +/- (83/6)i

OpenStudy (anonymous):

\[x = -\frac{1}{6} \pm \frac{83}{6}i\]

OpenStudy (anonymous):

Would have never got that ^

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

eri do you understand?

OpenStudy (anonymous):

Thank you so much!!!!!!

OpenStudy (anonymous):

You're welcome!!!